OpenStudy (anonymous):

A. write the expression 81^-3/4 in radical notation. B. evaluate the radical expression.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope im so lost

OpenStudy (skullpatrol):

Let's start with what a "negative" exponent means?

OpenStudy (skullpatrol):

$$\Huge a^{-n}= \frac{1}{a^n}$$

OpenStudy (skullpatrol):

Next, let's find out what a "fractional" exponent means?

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

\[\huge a ^{\frac{1}{n}}=\sqrt[n]{a}\]

OpenStudy (anonymous):

got the answer

OpenStudy (anonymous):

NO

OpenStudy (skullpatrol):

@sexybadass22 do you mean YOU got the answer?

OpenStudy (anonymous):

ok im lost

OpenStudy (anonymous):

NO. I AM CLUELESS!!!!

OpenStudy (anonymous):

ok me to

OpenStudy (anonymous):

ok im dam

OpenStudy (skullpatrol):

$$\huge 81^{-\frac{3}{4}}$$

OpenStudy (anonymous):

IT IS LINEAR PROGRAMING

OpenStudy (skullpatrol):